# Optimized Fish Locomotion using Design-by-Morphing and Bayesian Optimization

### v1

@@ 2.2.1 Modeling Undulatory Motion for Propulsion
where $y(x^{*},t)$ represents the lateral displacement of the body during oscillation, $2\pi/\lambda^{*}$ is the wave number, $f$ denotes the oscillation frequency, and $A(x^{*})$ is the swimming profile, non-dimensionalized by the total body length ($L$). The swimming profile $A(x^{*})$ plays a critical role in determining the efficiency and performance of propulsion. In this study, $A(x^{*})$ is optimized using a DbM technique. The optimized profile maximizes propulsion efficiency while considering constraints, such as stability and energy consumption.
@@
$m\frac{dU_{o}}{dt}=F_{T},\quad\frac{dx_{o}}{dt}=U_{o},$ (13a-b)
where $m=\rho V_{f}$ is the mass of the body, the density of the foil $\rho$ is regarded as a constant value equal to the density of the fluid in this work, $V_{f}$ is the volume of the foil, and $F_{T}$ represents the thrust force exerted by the fluid in the $x$-direction. The initial conditions are set as $U_{o}=0$ and $x_{o}=0$. The volume per unit depth of the NACA 0012 airfoil is considered as $0.08621\,\mbox{m}^{3}$, consistent with prior studies zhang2018effects .
The FSI problem is solved using the Hamming fourth-order predictor-corrector technique Viv_Mehmood2 , ensuring precise coupling between the fluid and structural dynamics. This computational framework enables the evaluation of the optimized profile $A(x^{*})$ by maximizing propulsive efficiency and minimizing energy dissipation.
@@ 2.2.2 Discretization Scheme
A pressure-correction algorithm is employed to solve the governing equations for fluid dynamics. Initially, the momentum equations are solved for velocity components in an uncoupled manner, neglecting the continuity constraint and omitting the pressure term. The resulting velocity field, referred to as the intermediate velocity, is then corrected to ensure mass conservation. This correction involves solving a pressure Poisson equation derived by imposing the continuity constraint on the velocity field. The computed pressure distribution is subsequently used to obtain a divergence-free velocity field by adjusting the intermediate velocity. This approach is consistent with the methodologies outlined by Kim and Moin kim1985application for Cartesian coordinates and Akhtar and Zang et al. ImranD ; cfd_Zang for curvilinear grids. In this study, an ‘O’-type body-fitted nearly orthogonal grid has been employed over the NACA-0012 airfoil (see Fig. 4), generated by numerically solving elliptic differential equations following the method of Ryskin and Leal ryskin1983orthogonal . A non-staggered grid arrangement is used, where the pressure and velocity components are computed at the cell center, while fluxes are evaluated at the midpoints of the cell faces.
The spatial discretization employs finite difference approximations. All spatial derivatives, except for convective terms, are approximated using a second-order accurate centered difference scheme. For the convective terms, a quadratic upwinding interpolation for convective kinematics (QUICK) scheme leonard1979stable is employed. This scheme utilizes a generic stencil based on positive and negative fluxes to enhance stability and accuracy. Temporal discretization is achieved using a semi-implicit scheme. The diagonal viscous terms are advanced using the Crank-Nicolson method for second-order accuracy, while all other terms are approximated using the explicit Adams-Bashforth method. For further details on the numerical schemes and boundary conditions, readers are referred to the following Refs. ImranD ; farooq2022nonlinear ; farooq2022numerical .
@@ 2.2.3 Re-meshing Strategy
The effectiveness of the ALE method largely depends on the re-meshing algorithm employed donea2017arbitrary . Various methods were proposed for re-meshing both structured and unstructured meshes batina1990unsteady ; degand2002three ; farhat1998torsional . In fluid dynamics, the most commonly used re-meshing techniques involve solving partial differential equations helenbrook2003mesh ; lohner1996improved ; dwight2009robust . However, these techniques often struggle to maintain mesh quality, especially when the body undergoes large rotational deformations.
Alternatively, a re-meshing technique developed by de2007mesh utilizes an interpolation approach based on the radial basis function (RBF) interpolation. This method is proven to be highly effective in maintaining mesh quality even under large translational and rotational displacements. Compared to other re-meshing strategies, the RBF method demonstrates superior efficiency and robustness bos2013radial . In our simulations, we employ this technique to determine the displacement of both boundary and interior nodes.
@@
where $N_{c}$ represents the number of control points, $\phi(\|\cdot\|)$ is the RBF as a function of the Euclidean norm. In this case, we use the global support thin-plate spline (TPS) radial basis function, $\phi(x)=x^{2}\log(x)$. The term $\mathbf{p}(\mathbf{x})$ is a minimal polynomial of degree one with coefficients $\mathbf{\beta}$, which must satisfy the following condition:
@@
Here the $\alpha_{i}$’s are the coefficients of the basis function that are determined under the condition that at each control point, $f(\mathbf{x})$ interpolates the given displacements, i.e., $f(\mathbf{xc})=\Delta\mathbf{xc}$, where $\Delta\mathbf{xc}$ is the prescribed displacement at the node $\mathbf{xc_{j}}$. The control points are selected from a subset of nodes on the airfoil boundary using a greedy algorithm rendall2009efficient .
@@
Further, to accomplish the propulsion of undulating bodies, we design a dynamical mesh shown in Figure 4 which indicates two zones, separated by an inner circle of radius 5c, only zone-1 is allowed to deform during the simulation at each time step based on the motion of the structural boundary nodes using the proposed RBF re-meshing algorithm, whereas the shape of the grid cells in zone-2 remains fixed. To propel the structure, all grid nodes are allowed to move ahead with the same velocity of $U_{o}$.
@@ 2.3 Bayesian Optimization
Optimizing expensive black-box functions, particularly when involving mixed-variable or high-dimensional design spaces, is an active area of research. BO is widely recognized as a highly sample-efficient strategy, as it achieves competitive performance with significantly fewer function evaluations compared to conventional optimization methods brochu2010tutorial . In the present study, the evaluation of each design case involves high-fidelity simulations of self-propelling undulatory foils, which are computationally expensive. Therefore, BO provides an appropriate choice of optimizer to ensure that the optimization framework remains computationally feasible.
Several engineering design problems utilized BO in mixed-variable or high-dimensional design spaces, including architected material design sheikh2022systematic ; vangelatos2021strength , hyperparameter tuning in machine learning oh2018bock , drug discovery korovina2020chembo , and fluid-structure optimization sheikh2023airfoil . In our case, the design space is parameterized by four morphing weights that define the swimming profile $A(x^{*})$ through design-by-morphing, along with two undulation parameters: frequency and wavelength. Thus, the total design space constitutes six continuous variables. For such expensive, noisy, and potentially multi-modal problems, local optima can trap conventional algorithms, while BO effectively balances exploration and exploitation.
To address this optimization problem, we employ the Mixed-Variable Multi-Objective Bayesian Optimization (MixMOBO) algorithm sheikh2022bayesian . This framework is capable of handling both continuous and categorical variables, and has demonstrated efficiency in optimizing noisy black-box functions with limited evaluation budgets. Although MixMOBO was previously applied in multi-objective settings for architected meta-materials and fluid-structure systems, here it is adopted in a single-objective optimization setting, with the goal of maximizing propulsive efficiency. MixMOBO combines multiple acquisition functions such as Expected Improvement, Probability of Improvement, Upper Confidence Bound, and Stochastic Monte Carlo under a hedging strategy, thereby improving robustness across diverse objective landscapes. Following best practices sheikh2022bayesian ; sheikh2022optimization , an initial sampling size of approximately 10–20% of the total budget is employed.
To validate the efficacy of MixMOBO for the present optimization problem, we also test it on a set of standard benchmark functions commonly used in the literature, namely the Spherical, Rastrigin, and Styblinski–Tang functions, see Fig. 5, along with a customized anisotropic function designed to mimic the characteristics of the swimming profile design space. The optimization performance is quantified using the normalized reward metric, defined as:
@@
where $f_{\text{opt}}$ is the optimum obtained by BO, $f_{\text{init}}$ is the optimum found by random search, and $f_{\text{global}}$ denotes the known global optimum.
Figure 5: Benchmarks for MixMOBO for four test functions.
3 Numerical Validation
[…]
Using grid G2, a time-step independence study is also performed by considering three levels of time-steps: TS1, 3000; TS2, 4000; and TS3, 5000 time-steps per oscillating cycle. As shown in Table 2, TS1 is chosen for time-marching in the simulations since the variation in $\bar{U}_{o}$ value with respect to different time-steps is minimal.
Time Steps
Grid Resolution per Cycle $\bar{U}_{o}$
@@ 3 Numerical Validation
Furthermore, we perform numerical simulations for flows around a self-propelling NACA 0012 airfoil under the same configuration as considered by Zhang et al. zhang2018effects and compare the obtained results with the data reported in Zhang et al. zhang2018effects . These simulations are performed for $\mbox{Re}=1000$ and undulatory frequencies varying from 0.4 Hz to 1.3 Hz. We compute the mean value of $U_{o}$, work consumption, and propulsive efficiency, and then plot these metrics in Fig. 6 versus the undulatory frequency, alongside the results of Zhang et al. zhang2018effects . Excellent agreement is found between our computed results and those reported in the literature.
Figure 6: Validation: comparison of (a) the mean forward velocity $\bar{U}_{o}$, (b) work consumption of the undulating foils, and (c) propulsive efficiency with the published data reported in zhang2018effects .
@@ 4 Results & Discussion
Our present work includes high-fidelity 2D simulations of self-propulsive undulating bodies for over a wide range of parameter space, involving four design space variables and two undulation parameters, namely frequency and wavelength. The specifications of the governing parameters, structural kinematics, design variables, and flow regimes are presented in Table 3. Furthermore, the subsequent sections present a detailed discussion of the findings from the optimization study.
@@
Before delving into the detailed findings of the objective optimization, it is essential to assess the performance of the commonly used carangiform and anguilliform modes swimming profiles in the frequency-wavelength space. To this end, high-fidelity simulations are conducted over the specified wavelength–frequency domain, resulting in 198 simulations, 99 for each profile. We compute the propulsive efficiency for both swimming modes. The results obtained from these simulations serve as benchmark references for evaluating the performance of optimized swimming profiles generated through the DbM-BO approach. By establishing a baseline using these well-studied natural modes, we can better quantify the improvements achieved by our optimization framework in terms of enhanced propulsive efficiency and reduced energy expenditure across the design space.
In Fig. 7, the propulsive efficiency contours across the frequency–wavelength domain are illustrated for both carangiform and anguilliform swimming modes at $\mbox{Re}=1000$. The maximum efficiency observed in both modes lies in the range of 17–22%, indicating the upper bound of performance under the given flow conditions and kinematic parameters. The efficiency of carangiform shows a clear trend, with the highest values observed in the lower left region of the domain, specifically for lower wavelength ($\lambda^{*}\approx 0.5$$0.7$) and lower to moderate frequency values ($f\approx 1.0$$2.5$). The peak efficiency reaches approximately 17.21%, indicating that carangiform swimming achieves optimal energy conversion under compact and relatively slower undulatory motions. In contrast to the carangiform mode, the anguilliform profile exhibits a broader region of elevated efficiency. The maximum efficiency reaches around 22%, primarily in the wavelength range ($\lambda^{*}\approx 1.1$$1.4$) and low frequencies ($f\approx 0.5$$1.5$). The distribution is more symmetric compared to carangiform, suggesting that anguilliform swimming maintains high efficiency over a wider range of undulatory wavelengths. These results serve as important baselines for evaluating the performance of swimming profiles optimized through the DbM-BO in the same frequency–wavelength domain.

### v2

@@ 2.2.1 Modeling Undulatory Motion for Propulsion
where $y(x^{*},t)$ denotes the lateral displacement of the body during oscillation, $2\pi/\lambda^{*}$ is the wave number, $f$ is the oscillation frequency, and $A(x^{*})$ represents the swimming profile non-dimensionalized by the total body arc-length $L$. The swimming profile $A(x^{*})$ plays a central role in determining the propulsive performance and efficiency and is optimized in this study using the DbM technique.
The coordinate $x^{*}$ is a normalized material coordinate that describes the position along the swimming body, independent of its physical length. The physical body length is determined by enforcing the preservation of the arc-length (and thus the conservation of volume). Accordingly, the physical $x$-coordinate is obtained by scaling the material coordinate as $x=\alpha x^{*}$, where $\alpha$ denotes the instantaneous physical length of the body. The head and tail positions are fixed at $x(0)=0$ and $x(1)=\alpha$, respectively.
To enforce a prescribed arc-length $L$, the scaling parameter $\alpha$ is obtained at each time step by solving the following nonlinear constraint:
$\int_{0}^{1}\sqrt{\alpha^{2}({dx^{*}})^{2}+({dy})^{2}}\,=L$ (13)
@@
$m\frac{dU_{o}}{dt}=F_{x},\quad\frac{dx_{o}}{dt}=U_{o},$ (14a-b)
where $m=\rho V_{f}$ is the mass of the body, the density of the foil $\rho$ is regarded as a constant value equal to the density of the fluid in this work, $V_{f}$ is the volume of the foil, and $F_{x}$ represents the force exerted by the fluid in the $x$-direction. The initial conditions are set as $U_{o}=0$ and $x_{o}=0$. The volume of the NACA 0012 airfoil is considered as $0.08621\,\mbox{m}^{3}$, consistent with prior studies (Zhang et al., 2018).
The FSI problem is solved using the Hamming fourth-order predictor-corrector technique (Mehmood et al., 2013), ensuring precise coupling between the fluid and structural dynamics. This computational framework enables the evaluation of the optimized profile $A(x^{*})$ by maximizing propulsive efficiency and minimizing energy dissipation.
@@ 2.2.2 Discretization Scheme
A pressure-correction algorithm is employed to solve the governing equations for fluid dynamics. Initially, the momentum equations are solved for velocity components in an uncoupled manner, neglecting the continuity constraint and omitting the pressure term. The resulting velocity field, referred to as the intermediate velocity, is then corrected to ensure mass conservation. This correction involves solving a pressure Poisson equation derived by imposing the continuity constraint on the velocity field. The computed pressure distribution is subsequently used to obtain a divergence-free velocity field by adjusting the intermediate velocity. This approach is consistent with the methodologies outlined by Kim and Moin (1985) for Cartesian coordinates and Akhtar (2008); Zang et al. (1994) for curvilinear grids. In this study, an ‘O’-type body-fitted nearly orthogonal grid has been employed over the NACA-0012 airfoil (see Fig. 4), generated by numerically solving elliptic differential equations following the method of Ryskin and Leal (1983). A non-staggered grid arrangement is used, where the pressure and velocity components are computed at the cell center, while fluxes are evaluated at the midpoints of the cell faces.
The spatial discretization employs finite difference approximations. All spatial derivatives, except for convective terms, are approximated using a second-order accurate centered difference scheme. For the convective terms, a quadratic upwinding interpolation for convective kinematics (QUICK) scheme (Leonard, 1979) is employed. This scheme utilizes a generic stencil based on positive and negative fluxes to enhance stability and accuracy. Temporal discretization is achieved using a semi-implicit scheme. The diagonal viscous terms are advanced using the Crank-Nicolson method for second-order accuracy, while all other terms are approximated using the explicit Adams-Bashforth method. For further details on the numerical schemes and boundary conditions, readers are referred to the following Refs. Akhtar (2008); Farooq et al. (2022b, a).
@@ 2.2.3 Re-meshing Strategy
The effectiveness of the ALE method largely depends on the re-meshing algorithm employed (Donea et al., 2017). Various methods were proposed for re-meshing both structured and unstructured meshes (Batina, 1990; Degand and Farhat, 2002; Farhat et al., 1998). In fluid dynamics, the most commonly used re-meshing techniques involve solving partial differential equations (Helenbrook, 2003; Löhner and Yang, 1996; Dwight, 2009). However, these techniques often struggle to maintain mesh quality, especially when the body undergoes large rotational deformations.
Alternatively, a re-meshing technique developed by de Boer et al. (2007) utilizes an interpolation approach based on the radial basis function (RBF) interpolation. This method is proven to be highly effective in maintaining mesh quality even under large translational and rotational displacements (Bos et al., 2013). Compared to other re-meshing strategies, the RBF method demonstrates superior efficiency and robustness. In our simulations, we employ this technique to determine the displacement of both boundary and interior nodes.
@@
where $N_{c}$ represents the number of control points, $\phi(\|\cdot\|)$ is the RBF as a function of the Euclidean norm. In this case, we use the global support thin-plate spline (TPS) radial basis function, $\phi(x)=x^{2}\log(x)$. The term $\mathbf{p}(\mathbf{x})$ is a minimal polynomial of degree one with coefficients vector $\boldsymbol{\beta}$, which must satisfy the following condition (Bos et al., 2013):
@@
Here the $\alpha_{i}$’s are the coefficients of the basis function that are determined under the condition that at each control point, $f(\mathbf{x})$ interpolates the given displacements, i.e., $f(\mathbf{xc})=\Delta\mathbf{xc}$, where $\Delta\mathbf{xc}$ is the prescribed displacement at the node $\mathbf{xc}$. The control points are selected from a subset of nodes on the airfoil boundary using a greedy algorithm (Rendall and Allen, 2009).
@@
To enable the self-propulsion of the undulating body, a dynamic mesh framework is employed, as illustrated in Fig. 4. The computational domain is divided into two zones, separated by an inner circular interface located at a radius of approximately $10L$. The two zones are constructed independently, not in a manner analogous to a sliding-mesh strategy. The orientations of the two zones shown in the figure are presented solely for illustrative purposes, to highlight the mesh deformation strategy. Only the inner region (Zone 1) is allowed to deform during the simulation, with grid motion prescribed at each time step based on the displacement of the structural boundary nodes using the proposed RBF-based re-meshing algorithm. In contrast, the grid topology and cell shapes in the outer region (Zone 2) remain fixed. To realize self-propulsion, all grid nodes are translated with a uniform streamwise velocity $U_{o}$, which is determined dynamically from the force balance on the body.
@@ 2.3 Bayesian Optimization
Optimizing expensive black-box functions, particularly when involving mixed-variable or high-dimensional design spaces, is an active area of research. BO is widely recognized as a highly sample-efficient strategy, as it achieves competitive performance with significantly fewer function evaluations compared to conventional optimization methods (Brochu et al., 2010). In the present study, the evaluation of each design case involves high-fidelity simulations of self-propelling undulatory foils, which are computationally expensive. Therefore, BO provides an appropriate choice of optimizer to ensure that the optimization framework remains computationally feasible.
Several engineering design problems utilized BO in mixed-variable or high-dimensional design spaces, including architected material design (Sheikh et al., 2022b; Vangelatos et al., 2021), hyperparameter tuning in machine learning (Oh et al., 2018), drug discovery (Korovina et al., 2020), and fluid-structure optimization (Sheikh et al., 2023). In our case, the design space is parameterized by four morphing weights that define the swimming profile $A(x^{*})$ through design-by-morphing, along with two undulation parameters: frequency and wavelength. Thus, the total design space constitutes six optimization input variables. For such expensive, noisy, and potentially multi-modal problems, local optima can trap conventional algorithms, while BO effectively balances exploration and exploitation.
To address this optimization problem, we employ the Mixed-Variable Multi-Objective Bayesian Optimization (MixMOBO) algorithm (Sheikh and Marcus, 2022). This framework is capable of handling both continuous and categorical variables, and has demonstrated efficiency in optimizing noisy black-box functions with limited evaluation budgets. In this study, MixMOBO is adopted in a single-objective discrete-parameter setting with the in-built HedgeMO strategy to ensure robustness, with the goal of maximizing propulsive efficiency. The design space is discretized to take advantage of MixMOBO’s efficiency when dealing with mixed-variables and to ensure that the algorithm does not get stuck during exploration in a high-frequency response region (Sheikh et al., 2022a). Following best practices (Sheikh and Marcus, 2022; Sheikh et al., 2022a), an initial sampling size of approximately 10–20% of the total budget is employed.
To validate the efficacy of MixMOBO for the present optimization problem, we also test it on a set of standard 6D benchmark functions commonly used in the literature, namely the Spherical, Rastrigin, Styblinski–Tang, and anisotropic Amalgamated test functions, with design space similar to our swimming profile optimization design space. The test functions definitions and settings are same as (Sheikh et al., 2022a). Each test function was run for 10 independent runs and performance was quantified using the normalized reward metric, defined as:
@@
where $f_{\text{opt}}$ is the optimum obtained by BO, $f_{\text{init}}$ is the optimum found by random search, and $f_{\text{global}}$ denotes the known global optimum. The results are demonstrated in Fig. 5 as the mean normalized reward with standard deviation over 10 runs. These results show MixMOBO can achieve consistent convergence within our prescribed computational budget of 150 epochs with 50 initial sample points for a design space of similar dimensionality.
Figure 5: MixMOBO performance benchmarking on four test functions.
3 Numerical Verification & Validation
This section details the verification of the numerical solver through systematic grid-independence and time-step refinement studies. In addition, the validation of the proposed methodology is performed by conducting simulations for both self-propelled and tethered swimming configurations and comparing the results with available reference data.
A detailed grid-independence study is performed using four grid resolutions: G1 with $250\times 200$ grid points, G2 with $374\times 300$ grid points, G3 with $512\times 400$ grid points, and G4 with $614\times 480$ grid points. Flow simulations are performed to compute the time-averaged drag coefficient (defined as $\bar{C}_{D}=\bar{F}_{D}/0.5\rho LU_{\infty}^{2}$) and the power ($\bar{P}=W/T$) for a tethered NACA 0012 airfoil undergoing carangiform undulation at an undulation frequency of 4 Hz and a non-dimensional wavelength of 0.5. The computed parameters are reported in Table 2. No significant changes in the computed parameters are observed when the resolution is increased beyond that of grid G3, indicating that the medium mesh size provides satisfactory results.
Using grid G3, a time-step independence study is also performed by considering three levels of time-steps: TS1, 3000; TS2, 4000; and TS3, 5000 time-steps per oscillating cycle. As shown in Table 2, TS2 is chosen for time-marching in the simulations since the variation in computed parameters with respect to different time-steps is minimal.
Grid Resolution Time Steps
$N_{x}\times N_{y}$ per Cycle $\bar{C}_{D}$ $\bar{P}$
@@ 3 Numerical Verification & Validation
[…]
For the validation of the numerical solver, we first compare our results for a tethered NACA 0012 airfoil undergoing carangiform undulation at $\mbox{Re}=500$ and $\lambda^{*}=1$ over a range of undulation frequencies from 1 to 4 Hz. Our results, in terms of the time history of drag coefficient, along with those reported by Khalid et al. (2016a), is plotted in Fig. 6, showing good agreement across all considered frequencies. In addition, simulations of a tethered NACA 0012 airfoil undergoing anguilliform undulation at $\mbox{Re}=1000$ are performed over a range of undulation frequencies and non-dimensional wavelengths. The resulting time-averaged power, as shown in Fig. 6, is compared with the results of Khalid et al. (2020c), confirming the accuracy of the solver in terms of power computation.
Furthermore, we perform numerical simulations of flow around a self-propelling NACA 0012 airfoil under the same configuration as considered by Zhang et al. (2018) and compare the obtained results with the data reported therein. These simulations are conducted at $\mbox{Re}=1000$ for undulatory frequencies ranging from 0.4 to 1.3 Hz. The mean forward velocity $U_{o}$, work consumption, and propulsive efficiency are computed and plotted as functions of the undulatory frequency in Fig. 7, together with the corresponding results of Zhang et al. (2018). Excellent agreement is observed between the present results and those available in the literature, further validating the coupling of the dynamical system with the numerical framework.
Figure 6: Validation of the present results: (a) time history of drag coefficients compared with the data of Khalid et al. (2016a), and (b) time-averaged power compared with the results of Khalid et al. (2020c).
Figure 7: Comparison of (a) the mean forward velocity $\bar{U}_{o}$, (b) work consumption of the undulating foils, and (c) propulsive efficiency with the published data reported in (Zhang et al., 2018) with $\lambda^{*}=1$.
@@ 4 Results & Discussion
Our present work includes 2D simulations of self-propulsive undulating bodies for over a wide range of parameter space, involving four design space variables and two undulation parameters, namely frequency and wavelength. The specifications of the governing parameters, structural kinematics, design variables, and flow regimes are presented in Table 3. All the cases of self-propulsion are simulated for 24 oscillating cycles to mitigate the transient effects. Furthermore, the subsequent sections present a detailed discussion of the findings from the optimization study.
@@
Before delving into the detailed findings of the objective optimization, it is essential to assess the performance of the commonly used carangiform and anguilliform swimming profiles in the frequency-wavelength space. To this end, high-fidelity simulations are conducted over the specified wavelength–frequency domain, resulting in 162 simulations, 81 for each profile. We compute the propulsive efficiency for both swimming modes. The results from these simulations serve as benchmarks for evaluating the performance of optimized swimming profiles generated by the DbM-BO approach. By establishing a baseline using these well-studied natural modes, we can better quantify the improvements achieved by our optimization framework in terms of enhanced propulsive efficiency across the design space.
In Fig. 8, the propulsive efficiency contours over the frequency–wavelength domain are shown for both carangiform and anguilliform swimming modes at $\mbox{Re}=1000$. For both modes, the maximum efficiency lies in the range of approximately $35\%$$42\%$, representing the upper bound of performance under the considered flow conditions and kinematic parameters. The carangiform mode exhibits a clear trend, with higher efficiencies concentrated in the lower-left region of the parameter space, corresponding to relatively shorter wavelengths ($\lambda^{*}\approx 0.5$$1.1$) and low to moderate undulation frequencies ($f\approx 0.2$$0.8$). The peak efficiency reaches about $42\%$, indicating that carangiform swimming achieves optimal energy conversion through compact body undulations at comparatively lower frequencies. In contrast, the anguilliform mode displays a broader region of elevated efficiency, primarily distributed along a vertical band spanning wavelengths $\lambda^{*}\approx 0.9$$1.5$ and frequencies $f\approx 0.2$$0.9$. This distribution is more symmetric than that observed for carangiform swimming, suggesting that anguilliform kinematics sustain high propulsive efficiency across a wider range of undulatory wavelengths and frequencies. These efficiency maps provide important baseline references for assessing the performance gains achieved by swimming profiles optimized using the DbM–BO framework within the same frequency–wavelength domain.